OpenStudy (anonymous):

haha for fun :P find the exact value of \(S=1+\frac{1}{2^6}+\frac{1}{3^6}+\frac{1}{4^6}+...\)

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ \dfrac{\pi^6}{945}\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

yes , but why :P

OpenStudy (mathmath333):

lol, wolfram gave it

OpenStudy (anonymous):

:P ohk lol show a method

OpenStudy (anonymous):

i only know 2 method, so looking for more excited answers

OpenStudy (anonymous):

@AkashdeepDeb do u memorize this ? what function for Taylor does this represent ! (sorry not convinced enough )

OpenStudy (kainui):

Well since \[\Large \frac{\pi^2}{6} = \frac{1}{1^2}+\frac{1}{2^2}+\frac{1}{3^2}+...\] and we can rewrite what you put as \[\Large \frac{1}{1^2} + \frac{1}{8^2}+\frac{1}{27^2}+\frac{1}{64^2}+...\] I can at least put the most amount of reasoning into an upper bound as pi^2/6 which at least has a method to it. =P

OpenStudy (anonymous):

i see well , same method used to solve first sum can solve the second one =) so yeah try , why not !

OpenStudy (anonymous):

ok what i know is using Fourier (but would be very messy) but the other easy method is Zeta function @ganeshie8 a question with zeta :)

OpenStudy (anonymous):

You could probably use some variation of Euler's proof for \(\sum\dfrac{1}{n^2}\) converging to \(\dfrac{\pi^2}{6}\). Where he used \(\dfrac{\sin x}{x}\), we (maybe) could get away with using \(\dfrac{\sin(x^3)}{x^3}\).

OpenStudy (anonymous):

exactly :)

OpenStudy (anonymous):

this is also euler approach he invented zeta

ganeshie8 (ganeshie8):

\[\zeta(2n) = \sum\limits_{i=1}^{\infty} \dfrac{1}{i^{2n}} = (-1)^{n+1} \frac{B_{2n} 2^{2n}}{2(2n)!} \pi^{2n}\] plugin \(n = 3\) ?

OpenStudy (anonymous):

Just curious, is there a way to write the infinite series as the limit of a finite series, and somehow translate that into a definite integral?

ganeshie8 (ganeshie8):

riemann sum @SithsAndGiggles ?

OpenStudy (anonymous):

Yes I was wondering if there's a way to rewrite as follows \[\sum_{n=1}^\infty \frac{1}{n^6}=\lim_{k\to\infty}\sum_{n=1}^kf(a+k\Delta x)\Delta x=\int_a^b f(x)\,dx\] but I can't seem to make it work...

ganeshie8 (ganeshie8):

\[\int_0^1 x^{-k}\,dx = \lim\limits_{n\to \infty} \sum\limits_{i=1}^n (\frac{i}{n})^{-k} \cdot \frac{1}{n}= \lim\limits_{n\to \infty} \frac{1}{n^{-k+1}}\sum\limits_{i=1}^n i^{-k} \] something like this ?

ganeshie8 (ganeshie8):

we know how to evaluate the left side definite integral but extracting that sum on right hand side looks trickyk.. xD

ganeshie8 (ganeshie8):

that doesn't look converging, we might have to tweak the bounds hmm

OpenStudy (anonymous):

yes exactly !

OpenStudy (kainui):

Here is an idea I had before but couldn't seem to take it very far, but I think I might have made some progress now, but I still need help. \[\Large \frac{\pi^2}{6} = \frac{1}{(1^1)^2}+\frac{1}{(2^1)^2}+ \frac{1}{(3^1)^2}+... \\ \Large \frac{\pi^4}{90} = \frac{1}{(1^2)^2}+\frac{1}{(2^2)^2}+ \frac{1}{(3^2)^2}+... \\ \Large \frac{\pi^6}{945} = \frac{1}{(1^3)^2}+\frac{1}{(2^3)^2}+ \frac{1}{(3^3)^2}+...\] But I was thinking, we can add up all the pi^2, pi^4, pi^6, etc terms to infinity and it would be equal to all of these geometric series if you look at them vertically going down the columns.\[\Large \sum_{k=1}^\infty \frac{1}{(n^k)^2}=\sum_{k=1}^\infty \left(\frac{1}{n^2} \right)^k=\frac{1}{n^2-1}\] So I was thinking maybe there's some way to set these equal to each other and solve some thing like this and derive them from something like this (despite being divergent) \[\Large \sum_{n=1}^\infty \frac{1}{n^2-1} = \sum_{n=1}^\infty \frac{\pi^{2n}}{constant?}\] So here's as far as I can go, I don't know if there's a simple way to represent these constants in the denominator, but maybe we can reverse this into some kind of derivation just like how fourier series sort of work.

OpenStudy (kainui):

Actually on wikipedia it shows it as \[\Large \zeta (2n) = (-1)^{n+1} \frac{B_{2n}(2 \pi)^{2n}}{2(2n)!}\] and B_n are the Bernoulli numbers (which I really don't know much about sadly, so if anyone can explain them to me that would be great, so I can rewrite my equation as: \[\Large \sum_{n=1}^\infty \frac{1}{n^2-1} = \sum_{n=1}^\infty (-1)^{n+1} \frac{B_{2n}(2 \pi)^{2n}}{2(2n)!}\] Which is kind of interesting, since it looks like we have the left side can factor as a difference of two squares, I don't know if this gets us anywhere closer to solving Ikram's problem the long way or something lol.

OpenStudy (kainui):

There are two ways I can see how to change the stuff I threw up here to make it convergent and hopefully I can find something interesting.

OpenStudy (kainui):

Ok so instead of adding them all up divergently, let's just add zeta(2)-zeta(4)+zeta(6)-... etc and change the geometric series accordingly to have a negative sign in it! That will make it converge and change the expression to be: \[\Large \sum_{n=1}^\infty \frac{1}{n^2+1} = \sum_{n=1}^\infty \frac{B_{2n}(2 \pi)^{2n}}{2(2n)!}\] My first instinct is to associate the integral on the left with this integral: \[\Large \int\limits_0^\infty \frac{dx}{x^2+1}=\frac{\pi}{2}\] But then I also noticed that this summation on the left is really the shifted basel problem, which is strange, maybe there's an easy way to solve these geometrically afterall? I am sort of starting to believe that there is some geometric meaning by having pi in there. Maybe you find this interesting as well?

ganeshie8 (ganeshie8):

\[\sum\limits_{k=1}^{n} k^m = \frac{1}{m+1} \sum \limits_{k=0}^m \binom{m+1}{k} \color{Red}{B_k} n^{m+1-k}\] we can generate all bernouli numbers by using that definition http://en.wikipedia.org/wiki/Bernoulli_number

ganeshie8 (ganeshie8):

idk much about zeta function but i have seen bernouli numbers in number theory before @Marki promised she will teach me zeta someday... fascinating stuff xD that integral reminds me of euler reflection formula : \[\int\limits_0^{\infty} \dfrac{dx}{x^n+1} = \dfrac{\Gamma(1-\frac{1}{n}) \Gamma(\frac{1}{n})}{n} = \dfrac{\pi}{n\sin(\frac{\pi}{n})}\] not sure if there will be some useful connection with the discrete sum :/

OpenStudy (anonymous):

haha well u two said 2 imp. things 1- it really has something to do with solving Bessel equation 2-Euler derived it , no doubt ur gonna smell Euler :) and @ganeshie8 i like this way in teaching step by step i believe now ur be able to find zeta(2n) for solving series of this form :) get ready for next step

OpenStudy (anonymous):

@Marki Bessel equation, or Basel problem?

OpenStudy (kainui):

Basel, Bessel, Bernoulli... What's the difference? =P

OpenStudy (anonymous):

i think problem thx for correcting me @SithsAndGiggles

OpenStudy (anonymous):

Idk maybe there's some crazy connection to the Bessel equation too...

OpenStudy (anonymous):

maybe , i'll see after exams

ganeshie8 (ganeshie8):

good luck wid exams @Marki

OpenStudy (anonymous):

:)